OpenStudy (anonymous):

If 3sin^2A=2.25 and A is an acute angle,then find A. I started with this problem but am always getting stuck at a point!!

OpenStudy (mathstudent55):

3sin^2A = 2.25 sin^2A = 0.75

OpenStudy (anonymous):

ya! then i took the square root of 0.75!! then?

OpenStudy (mathstudent55):

sinA = sqrt(0.75) = sqrt(3/4) = sqrt(3)/2

OpenStudy (mathstudent55):

Don't you recognize sqrt(3)/2 from a previous problem?

OpenStudy (anonymous):

Ya!! now i will be able proceed thanks!!!!

OpenStudy (mathstudent55):

wlcm